OpenStudy (anonymous):

Matthew is 3 times as old as Jenny. In 7 years, he will be twice as old as she will be then. How old is each now?

OpenStudy (shane_b):

Let M = Matthew and J = Jenny: M=3J M-3J=0 [1st equation] M+7=2(J+7)= M=2J+14-7 M=2J+7 M-2J=7 [2nd equation] Solve the system of both equations to get M = 21 and J = 7.

OpenStudy (anonymous):

Thank you :)